OpenStudy (anonymous):

Trying to figure out how to find the length of a polar curve eq - r=7sin(theta + Pi/4). I do not know how to do this HELP

OpenStudy (bahrom7893):

what is the equation? r = 7Sin(theta+Pi/4) or - r = 7Sin(theta+Pi/4) ?

OpenStudy (anonymous):

r=

OpenStudy (bahrom7893):

okay so..

OpenStudy (bahrom7893):

\[Length = \int\limits_{a}^{b} \sqrt{r^2+(dr/d \theta)^2}d \theta\]

OpenStudy (bahrom7893):

Now the hard part here is, finding the limits.. i forgot how to do that, let me graph this thing, see what it looks like..

OpenStudy (bahrom7893):

still tryin to figure out limits...

OpenStudy (bahrom7893):

do you know how to find limits? I've always been bad at polar stuff...

OpenStudy (anonymous):

no - this problem came from my College algebra/trig 400 class. Not doing integration yet.

OpenStudy (bahrom7893):

I think i know the limits, but sorry I can't solve this without integrating, never learned any other way of finding arclength..=(